### v1

# Alon’s transmitting problem and multicolor Beck–Spencer Lemma

Abstract.
The Hamming graph $H(n,q)$ is defined on the vertex set $\{1,2,\ldots,q\}^{n}$ and two vertices are adjacent if and only if they differ in precisely one coordinate. Alon [1] proved that for any sequence $v_{1},\ldots,v_{b}$ of $b=\lceil\frac{n}{2}\rceil$ vertices of $H(n,2)$, there is a vertex whose distance from $v_{i}$ is at least $b-i+1$ for all $1\leq i\leq b$. In this note, we prove that for any $q\geq 3$ and any sequence $v_{1},\ldots,v_{b}$ of $b=\lfloor(1-\frac{1}{q})n\rfloor$ vertices of $H(n,q)$, there is a vertex whose distance from $v_{i}$ is at least $b-i+1$ for all $1\leq i\leq b$.
Alon used the Beck–Spencer Lemma [3] which, in turn, was based on the floating variable method introduced by Beck and Fiala [2] who studied combinatorial discrepancies. For our proof, we extend the Beck–Spencer Lemma by using a multicolor version of the floating variable method due to Doerr and Srivastav [5].
1. Introduction
Alon posed a transmitting problem in [1] and obtained an optimal solution. This problem can be described in terms of graph burning number (see, e.g., [4]). Let $G$ be a finite graph with the vertex set $V$. For vertices $u,v\in V$ let $d(u,v)$ denote the distance between $u$ and $v$. For a non-negative integer $k$, let $\Gamma_{k}(v)$ denote the $k$-neighbors of $v$, that is, the set of vertices $u\in V$ such that $d(u,v)\leq k$. For $v_{1},v_{2},\ldots,v_{b}\in V$ we say that $(v_{1},v_{2},\ldots,v_{b})$ is a burning sequence of length $b$ if $\Gamma_{b-1}(v_{1})\cup\Gamma_{b-2}(v_{2})\cup\cdots\cup\Gamma_{0}(v_{b})=V$. The burning number of $G$, denoted by $b(G)$, is defined to be the minimum length of the burning sequences. We can think of the vertices as processors. Suppose that there is a sender outside the graph, and it sends a message to $v_{i}$ at round $i$. Each processor that receives a message at round $j$ transmits the message to all its neighbors at round $j+1$. Then $b(G)$ is the minimum number of rounds in which all the processors share the message.
Theorem 1 (Alon [1]).
Let $G$ be the $n$-dimensional cube. Then $b(G)=\lceil\frac{n}{2}\rceil+1$.
For the $n$-dimensional cube, let $v_{1}$ be an arbitrary vertex, and let $v_{2}$ be the antipodal vertex. Then it is easy to check that $\Gamma_{b-1}(v_{1})\cup\Gamma_{b-2}(v_{2})$ covers all the vertices where $b=\lceil\frac{n}{2}\rceil+1$. This means that the burning number is at most $b$. The more difficult part is to show that no matter how $b-1$ vertices are chosen, they cannot be a burning sequence, that is, after $b-1$ rounds there is still some vertex that has not received the message yet.
In this note, we extend the above result to Hamming graphs. For positive integers $n,q$, let $[q]:=\{1,2,\ldots,q\}$ and let $[q]^{n}$ denote the set of $n$-tuple of elements of $[q]$. The Hamming distance between $u,v\in[q]^{n}$ is defined to be the number of entries (coordinates) that they differ. The Hamming graph $H(n,q)$ has the vertex set $V=[q]^{n}$, and two vertices are adjacent if they have Hamming distance one. Note that $H(n,2)$ is (isomorphic to) the $n$-dimensional cube. Our main result is the following.
Theorem 2.
Let $q\geq 3$ be an integer, and let $G=H(n,q)$ be the Hamming graph. Then
$\left\lfloor\left(1-\tfrac{1}{q}\right)n\right\rfloor+1\leq b(G)\leq\left\lfloor\left(1-\tfrac{1}{q}\right)n+\tfrac{q+1}{2}\right\rfloor.$
The upper bound in Theorem 2 is easily verified by construction. Indeed, the sequence $v_{i}=(i,i,\ldots,i)$ for $i=1,2,\ldots,q$ (and any $v_{i}$ for $i>q$) works as a corresponding burning sequence (see [6] for more details). This construction is valid for $q=2$ as well, and in this case the upper bound coincides with the correct value of $b(H(n,2))$ given by Theorem 1.
To show the lower bound in Theorem 1, Alon used the Beck–Spencer lemma.
Lemma 1 (Beck–Spencer [3]).
For $1\leq i\leq n$, let ${\bm{a}}_{i}\in\{-1,1\}^{n}$ be given. Then there exists ${\bm{x}}\in\{-1,1\}^{n}$ such that the standard inner product satisfies $|{\bm{a}}_{i}\cdot{\bm{x}}|<2i$ for all $1\leq i\leq n$.
We can think of $\{-1,1\}^{n}$ as the vertex set of the $n$-dimensional cube, and it follows that ${\bm{a}}\cdot{\bm{x}}=n-2d({\bm{a}},{\bm{x}})$ as we will see in the next section. Then Lemma 1 is restated as follows.
Lemma 2 (Lemma 1 restated).
For $1\leq i\leq n$, let $v_{i}$ be a given vertex of $H(n,2)$. Then there exists a vertex $w$ such that $|n-2d(v_{i},w)|<2i$ for all $1\leq i\leq n$.
For our proof of the lower bound in Theorem 2, we need a multicolor version of the Beck–Spencer lemma.
Lemma 3.
Let $q\geq 3$. For $1\leq i\leq n$, let $v_{i}$ be a given vertex of $H(n,q)$. Then there exists a vertex $w$ such that $|(1-\frac{1}{q})n-d(v_{i},w)|<i$ for all $1\leq i\leq n$.
To prove Lemma 1 Beck and Spencer used the so-called floating variable method based on [2] where Beck and Fiala studied combinatorial discrepancies. Then Doerr and Srivastav in [5] extended the results in [2] (and many other related results concerning discrepancies) to multicolor settings. We utilize their ideas on vector-coloring to prove Lemma 3, c.f. the proof of Theorem 4.2 in [5]. So we explain a multicolor version of the floating variable method in the next section.
It follows from Lemma 1 that $b(H(n,2))\geq\lfloor\frac{n}{2}\rfloor+1$. However, this is not the sharp bound if $n$ is odd. Alon made one more twist to improve the bound in [1]. The author was not able to find an appropriate extension of this tricky part for $H(n,q)$ ($q\geq 3$), and this suggests that the lower bound in Theorem 2 could be improved.
Conjecture.
$b(H(n,q))=\left\lfloor\left(1-\tfrac{1}{q}\right)n+\tfrac{q+1}{2}\right\rfloor$.
2. Proofs
We want to restate Lemma 3 in the form of Lemma 1. For this, we need to assign a vector to a vertex in $H(n,q)$ so that the vector behaves nicely with respect to the inner product and the Hamming distance. To this end, we use vector-coloring introduced by Doerr and Srivastav in [5]. For a $qn$-dimensional vector ${\bm{z}}=(z_{1},\ldots,z_{qn})\in{\mathbb{R}}^{qn}$, let $\bar{z}_{j}:=(z_{q(j-1)+1},z_{q(j-1)+2},\ldots,z_{qj})\in{\mathbb{R}}^{q}$, and we also write ${\bm{z}}=(\bar{z}_{1},\ldots,\bar{z}_{n})$. (We write $\bar{z}$ to indicate the vector is $q$-dimensional.) For each color $i\in[q]$, we assign a vector $\bar{c}_{i}=-\frac{1}{q}\bar{1}+\bar{e}_{i}\in{\mathbb{R}}^{q}$, that is, $\bar{c}_{i}$ has $1-\frac{1}{q}$ at the $i$-th entry, and $-\frac{1}{q}$ at the other $q-1$ entries. For example, if $q=3$ then $\bar{c}_{1}=(\frac{2}{3},-\frac{1}{3},-\frac{1}{3})$, $\bar{c}_{2}=(-\frac{1}{3},\frac{2}{3},-\frac{1}{3})$, $\bar{c}_{3}=(-\frac{1}{3},-\frac{1}{3},\frac{2}{3})$. By definition the sum of all entries of $c_{i}$ equals 00. Let $Q$ be the set of color vectors, that is,
$Q=\{\bar{c}_{1},\bar{c}_{2},\ldots,\bar{c}_{q}\}.$
For each vertex $v=(\nu_{1},\ldots,\nu_{n})\in[q]^{n}$ of $H(n,q)$, we assign a vector ${\bm{z}}=(\bar{z}_{1},\ldots,\bar{z}_{n})\in{\mathbb{R}}^{qn}$ by
(1) $\bar{z}_{i}:=\bar{c}_{\nu_{i}}$
for $1\leq i\leq n$. In this case we understand ${\bm{z}}\in Q^{n}$. Define a bijection $\varphi:[q]^{n}\to Q^{n}$ by (1) and
(2) $\varphi(v):={\bm{z}}.$
Claim 1.
Let $v,w\in[q]^{n}$, and ${\bm{a}}=\varphi(v),{\bm{x}}=\varphi(w)$. Then ${\bm{a}}\cdot{\bm{x}}=(1-\frac{1}{q})n-d(v,w)$.
Proof.
Let $s:=\#\{i\in[n]:\bar{a}_{i}\neq\bar{x}_{i}\}=d(v,w)$. It follows from (3) that ${\bm{a}}\cdot{\bm{x}}=(1-\tfrac{1}{q})(n-s)-\tfrac{1}{q}s=(1-\tfrac{1}{q})n-d(v,w)$. ∎
Now we can restate Lemma 3.
Lemma 4 (Lemma 3 restated).
For $1\leq i\leq n$, let ${\bm{a}}_{i}\in Q^{n}$ be given. Then there exists ${\bm{x}}\in Q^{n}$ such that $|{\bm{a}}_{i}\cdot{\bm{x}}|<i$ for all $1\leq i\leq n$.
Proof.
Let $\tilde{Q}\subset{\mathbb{R}}^{q}$ denote the convex hull of $Q$, that is,
$\tilde{Q}:=\left\{\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}:0\leq\lambda_{i}\leq 1\,(i\in[q]),\,\sum_{i=1}^{q}\lambda_{i}=1\right\}.$
Claim 2.
For $\bar{a},\bar{x}\in Q$ and $\bar{y}\in\tilde{Q}$, we have the following.
  • $\bar{a}\cdot\bar{x}\in\{-\frac{1}{q},1-\frac{1}{q}\}$.
  • $-\frac{1}{q}\leq\bar{a}\cdot\bar{y}\leq 1-\frac{1}{q}$. If $\bar{a}\cdot\bar{y}=1-\frac{1}{q}$, then $\bar{y}=\bar{a}$.
  • $|\bar{a}\cdot\bar{x}-\bar{a}\cdot\bar{y}|\leq 1$. Moreover, if equality holds, then (i) $\bar{y}\in Q$ or (ii) there exists $i$ such that $\bar{a}=\bar{x}=\bar{c}_{i}$ and $i$-th entry of $\bar{y}$ is $-\frac{1}{q}$.
Proof.
Without loss of generality we may assume that $\bar{a}=\bar{c}_{1}$.
The first item follows from
(3) $\bar{c}_{i}\cdot\bar{c}_{j}=\begin{cases}1-\frac{1}{q}&\text{if $i=j$},\\ -\frac{1}{q}&\text{if $i\neq j$}.\end{cases}$
For the second item, let $\bar{y}=\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}\in\tilde{Q}$. Noting that $\sum_{i=2}^{q}\lambda_{i}=1-\lambda_{1}$, we have
(4) $\bar{a}\cdot\bar{y}=\bar{c}_{1}\cdot\sum_{i=1}^{q}\lambda_{i}\bar{c}_{i}=(1-\tfrac{1}{q})\lambda_{1}-\tfrac{1}{q}(1-\lambda_{1})=\lambda_{1}-\tfrac{1}{q}.$
Then $-\frac{1}{q}\leq\bar{a}\cdot\bar{y}\leq 1-\frac{1}{q}$ follows from $0\leq\lambda_{1}\leq 1$. If $\bar{a}\cdot\bar{y}=1-\frac{1}{q}$, then $\lambda_{1}=1$, that is, $\bar{y}=\bar{c}_{1}$.
By the first and second items we have $|\bar{a}\cdot\bar{x}-\bar{a}\cdot\bar{y}|\leq 1$. If equality holds, then $(\bar{a}\cdot\bar{x},\bar{a}\cdot\bar{y})=(-\frac{1}{q},1-\frac{1}{q})$ or $(1-\frac{1}{q},-\frac{1}{q})$. For the former case, we have $\bar{y}=\bar{a}\in Q$ by the second item. For the latter case, we have $\lambda_{1}=0$ by (4), and the first entry of $\bar{y}$ is $\lambda_{1}(1-\frac{1}{q})+\sum_{i=2}^{q}\lambda_{i}(-\frac{1}{q})=0-\frac{1}{q}(1-\lambda_{1})=-\frac{1}{q}$. ∎
Let ${\bm{x}}=(x_{1},\ldots,x_{qn})=(\bar{x}_{1},\ldots,\bar{x}_{n})\in{\mathbb{R}}^{qn}$ be a variable vector. We describe an algorithm updating ${\bm{x}}$ step by step in such a way that at each step ${\bm{x}}\in\tilde{Q}^{n}$, and in the end ${\bm{x}}\in Q^{n}$. We call a variable $x_{k}$ floating if $x_{k}\in\{-\frac{1}{q},1-\frac{1}{q}\}$, and fixed otherwise. Once a variable is fixed, then it stays the same value and it is treated as a constant.
We consider the following three conditions.
  • (C1)
    ${\bm{a}}_{i}\cdot{\bm{x}}=0$,
  • (C2)
    $x_{q(j-1)+1}+x_{q(j-1)+2}+\cdots+x_{qj}=0$,
  • (C3)
    ${\bm{x}}\in\tilde{Q}^{n}$.
We will solve a system of equations satisfying these conditions according to the following procedure.
Step(0) :
Consider the following system Eq(0) of equations:
    (C1) for $1\leq i\leq n$,
    (C2) for $1\leq j\leq n$.
Then ${\bm{x}}_{0}:={\mathbf{0}}$ is a solution to Eq(0), and it satisfies (C3) as well.
Step(1) :
Consider the following system Eq(1) of equations:
    (C1) for $11\leq i\leq n-$ 1,
    (C2) for $1\leq j\leq n$.
Eq(1) has $qn$ variables and $12n-$ 1 equations. Thus we have a non-trivial solution ${\bm{y}}\neq{\mathbf{0}}$. For every $\lambda\geq 0$, $\lambda{\bm{y}}$ also satisfies Eq(1). Moreover, if $\lambda=0$, then it satisfies (C3) as well.
Increase $\lambda$ continuously starting from 00. Then at some point at least one of the entries of $\lambda{\bm{y}}$ becomes $-\frac{1}{q}$ or $1-\frac{1}{q}$ for the first time. Using this $\lambda>0$, let ${\bm{x}}_{1}:={\bm{x}}_{0}+\lambda{\bm{y}}$. Then ${\bm{x}}_{1}$ satisfies Eq(1) and (C3). This ${\bm{x}}_{1}$ is not yet determined and may be updated. Let $s=1$ and go to Step(s) below.
Step(s) :
We have a temporary vector ${\bm{x}}_{s}=(\bar{x}_{1},\ldots,\bar{x}_{n})$ which satisfies (C1) for $1\leq i\leq n-s$, (C2) for $1\leq j\leq n$, and (C3). For $r\in\{0,1,\ldots,q\}$ let
$F_{r}({\bm{x}}_{s}):=\{j\in[n]:\text{the number of fixed entries in $\bar{x}_{j}$ is $r$}\}.$
Note that $F_{q-1}({\bm{x}}_{s})=\emptyset$ because if $q-1$ entries in $\bar{x}_{j}$ are fixed, then the remaining entry is necessarily in $\{-\frac{1}{q},1-\frac{1}{q}\}$ by (C2) and (C3).
If $|F_{q}({\bm{x}}_{s})|\geq s$, then we determine ${\bm{x}}_{s}$ and complete Step($s$). Moreover, if $s=n$, then exit the procedure, otherwise let ${\bm{x}}_{s+1}:={\bm{x}}_{s}$ and proceed to Step($s+1$). Note that ${\bm{x}}_{s+1}$ satisfies (C1) for $1\leq i\leq n-(s+1)$, (C2) for $1\leq j\leq n$, and (C3).
From now on, we deal with the case $|F_{q}({\bm{x}}_{s})|\leq s-1$. Since $|F_{q}({\bm{x}}_{s})|$ is non-decreasing in $s$, and $|F_{q}({\bm{x}}_{s-1})|\geq s-1$ holds when ${\bm{x}}_{s-1}$ is determined, we may assume that $|F_{q}({\bm{x}}_{s})|=s-1$. Let $f_{r}=|F_{r}({\bm{x}}_{s})|$. Then the number of fixed entries in ${\bm{x}}_{s}$ is $\sum_{r=1}^{q}rf_{r}$, and that of floating entries is $qn-\sum_{r=1}^{q}rf_{r}$.
We want to consider a system of equations consisting of only floating entries of ${\bm{x}}_{s}$. To this end, let (C1) and (C2) be equations obtained form (C1) and (C2), respectively, by replacing the coefficients corresponding to the fixed entries with zeros. Consider the following system Eq($s$) of equations:
    (C1) for $11\leq i\leq n-s=n-f_{q}-$ 1,
    (C2) for $j\in F_{0}({\bm{x}}_{s})\cup F_{1}({\bm{x}}_{s})\cup\cdots\cup F_{q-2}({\bm{x}}_{s})$.
Now we have $qn-\sum_{r=1}^{q}rf_{r}$ variables, and $(n-f_{q}-1)+(n-f_{q})=2n-2f_{q}-1$ equations. Thus the dimension of solution space is $(q-2)n-\sum_{r=1}^{q}rf_{r}+2f_{q}+1\geq 2f_{q}+1\geq 1$, and we have a non-trivial solution ${\bm{y}}\neq{\mathbf{0}}$. For every $\lambda\geq 0$, $\lambda{\bm{y}}$ also satisfies Eq($q$). We extend ${\bm{y}}\in{\mathbb{R}}^{[qn]\setminus F_{q}({\bm{x}}_{s})}$ to ${\bm{y}}\in{\mathbb{R}}^{qn}$ by inserting 00 at the positions in $F_{q}({\bm{x}}_{s})$. By continuously increasing $\lambda$ from 00, we get $\lambda{\bm{y}}$ such that one of the entries becomes $-\frac{1}{q}$ or $1-\frac{1}{q}$ for the first time. Using this $\lambda>0$, let ${\bm{x}}:={\bm{x}}_{s}+\lambda{\bm{y}}$. Then ${\bm{x}}$ satisfies (C1) for $1\leq i\leq n-s$, (C2) for $1\leq j\leq n$, and (C3). Update ${\bm{x}}_{s}:={\bm{x}}$, and return to the beginning of Step(s).
Completion of procedure :
When Step(n) is completed, we obtain ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$. By construction, we have the following conditions.
    $|F_{q}({\bm{x}}_{s})|\geq s$ and ${\bm{x}}_{s}\in\tilde{Q}^{n}$ for $1\leq s<n$.
    $|F_{q}({\bm{x}}_{n})|=n$, that is, ${\bm{x}}_{n}\in Q^{n}$.
    ${\bm{a}}_{i}\cdot{\bm{x}}_{n-i}=0$ for $1\leq i\leq n$. (Indeed ${\bm{a}}_{i}\cdot{\bm{x}}_{s}=0$ for $1\leq s<n$, $1\leq i\leq n-s$.)
By running this algorithm, determine ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$, and let ${\bm{x}}:={\bm{x}}_{n}=(\bar{x}_{1},\ldots,\bar{x}_{n})\in Q^{n}$.
We have ${\bm{a}}_{1}\cdot{\bm{x}}_{n-1}=0$. If ${\bm{x}}={\bm{x}}_{n-1}$, then $|{\bm{a}}_{1}\cdot{\bm{x}}|=0$. If ${\bm{x}}\neq{\bm{x}}_{n-1}$, then $|F_{q}({\bm{x}}_{n-1})|=n-1$, and there exists precisely one $j$ such that $j\not\in F_{q}({\bm{x}}_{n-1})$. Then, writing ${\bm{a}}_{1}=(\bar{a}_{1},\ldots,\bar{a}_{n})$ and ${\bm{x}}_{n-1}=(\bar{y}_{1},\ldots,\bar{y}_{n})$, we have $\bar{a}_{j},\bar{x}_{j}\in Q$, $\bar{y}_{j}\in\tilde{Q}\setminus Q$, and by Claim 2,
$|{\bm{a}}_{1}\cdot{\bm{x}}|=|{\bm{a}}_{1}\cdot{\bm{x}}-{\bm{a}}_{1}\cdot{\bm{x}}_{n-1}|=|\bar{a}_{j}\cdot\bar{x}_{j}-\bar{a}_{j}\cdot\bar{y}_{j}|\leq 1.$
Moreover, if $|\bar{a}_{j}\cdot\bar{x}_{j}-\bar{a}_{j}\cdot\bar{y}_{j}|=1$, then there exists $i$ such that $\bar{a}_{j}=\bar{x}_{j}=\bar{c}_{i}$, and the $i$-th entry of $\bar{y}_{j}$ is $-\frac{1}{q}$. But then the $i$-th entry of $\bar{y}_{j}$ is fixed and remains unchanged thereafter. This contradicts the fact that the $i$-th entry of $\bar{x}_{j}=\bar{c}_{i}$ is $1-\frac{1}{q}$. Consequently we have $|{\bm{a}}_{1}\cdot{\bm{x}}|=|\bar{a}_{j}\cdot\bar{x}_{j}-\bar{a}_{j}\cdot\bar{y}_{j}|<1$.
Similarly, for $1\leq i\leq n$, we have ${\bm{a}}_{i}\cdot{\bm{x}}_{n-i}=0$ and $F_{q}({\bm{x}}_{n-i})\geq n-i$. Thus, letting $J=[n]\setminus F_{q}({\bm{x}}_{n-i})$, we have $|J|\leq i$ and
$|{\bm{a}}_{i}\cdot{\bm{x}}|=|{\bm{a}}_{i}\cdot{\bm{x}}-{\bm{a}}_{i}\cdot{\bm{x}}_{n-i}|=\left|\sum_{j\in J}(\bar{a}_{j}\cdot\bar{x}_{j}-\bar{a}_{j}\cdot\bar{y}_{j})\right|\leq\sum_{j\in J}\left|\bar{a}_{j}\cdot\bar{x}_{j}-\bar{a}_{j}\cdot\bar{y}_{j}\right|<i,$
where we write ${\bm{a}}_{i}=(\bar{a}_{1},\ldots,\bar{a}_{n})$ and ${\bm{x}}_{n-i}=(\bar{y}_{1},\ldots,\bar{y}_{n})$. This completes the proof of Lemma 4. ∎
Proof of Theorem 2.
The upper bound is proved in [6]. Here we prove the lower bound. Let $m:=\lfloor(1-\tfrac{1}{q})n\rfloor$. For arbitrary $m$ vertices $v_{1},v_{2},\ldots,v_{n}\in[q]^{n}$, we show that there exists a vertex $w\in[q]^{n}$ such that $d(v_{i},w)\geq m+1-i$ for all $1\leq i\leq m$.
Recall the definition of $\varphi$ from (2). Let ${\bm{a}}_{i}=\varphi(v_{i})\in Q^{n}$ for $1\leq i\leq n$. By Lemma 4 and Claim 1, we get ${\bm{x}}\in Q^{n}$ such that
$|{\bm{a}}_{i}\cdot{\bm{x}}|=\left|(1-\tfrac{1}{q})n-d(v_{i},w)\right|<i$
for all $1\leq i\leq n$, where $w=\varphi^{-1}({\bm{x}})$. Thus we have
$d(v_{i},w)>(1-\tfrac{1}{q})n-i$
for all $1\leq i\leq n$. Let $n=qk+r$, $0\leq r<q$. Then,
$(1-\tfrac{1}{q})n=(q-1)k+r-\tfrac{r}{q}.$
This together with the fact that the distance is an integer yields
$d(v_{i},w)$$\geq\begin{cases}(q-1)k-i+1&\text{if $r=0$},\\ (q-1)k+r-i&\text{if $1\leq r<q$}\end{cases}$
$=m+1-i.$
Thus $w\not\in\Gamma_{m-1}(v_{1})\cup\cdots\cup\Gamma_{0}(v_{m})$, which means that $b(H,q)\geq m+1$. ∎
3. Concluding remarks
For the proof of Theorem 1, Alon observed the following. In Lemma 1 we have $|{\bm{a}}_{i}\cdot{\bm{x}}|=|{\bm{a}}_{i}\cdot(-{\bm{x}})|$, and so the inequality on $d(v_{i},w)$ in Lemma 2 also holds if we replace $v_{i}$ with its antipodal vertex. It would be nice if a similar extension was possible for Lemma 3, which would improve Theorem 2. For simplicity, let $q=3$ and $n=3k+1$. For $v=(x_{1},\ldots,x_{n})\in\{0,1,2\}^{n}$, let $v^{\prime}=(x_{1}+1,\ldots,x_{n}+1)$ and $v^{\prime\prime}=(x_{1}+2,\ldots,x_{n}+2)$, where addition is done in modulo 3. For vertices $u$ and $v$, let $f(u,v):=|\frac{2}{3}(3k+1)-d(u,v)|$, and $g(u,v):=\max\{f(u,v),f(u,v^{\prime}),f(u,v^{\prime\prime})\}$.
Problem.
For $11\leq i\leq 3k+$ 1, let $u_{i}$ be a given vertex of $H(3k+1,3)$. Then is it true that there exists a vertex $w$ such that $g(u_{i},w)<i$ for all $11\leq i\leq 3k+$ 1?
If this is true, then it implies that $b(H(3k+1,3))=2k+2$. To see this, assume that the answer to the problem is affirmative, and let vertices $v_{1},v_{2},\ldots,v_{2k+1}$ be given in $H(3k+1,3)$. As in [1], we apply the existence of $w$ to $u_{1}:=v_{2},u_{2}:=v_{3},\ldots,u_{2k}:=v_{2k+1}$. Then we have $g(v_{i},w)<i-1$ for $i=2,3,\ldots,2k+1$. Thus all of $w,w^{\prime}$ and $w^{\prime\prime}$ have distance at least $2k+2-i$ from $v_{i}$ for all $12\leq i\leq 2k+$ 1. Finally we use the fact that one of $w,w^{\prime}$, and $w^{\prime\prime}$ has distance at least $12k+$ 1 from $v_{1}$ because $d(v_{1},w)+d(v_{1},w^{\prime})+d(v_{1},w^{\prime\prime})=2(3k+1)$. Consequently, there is $w^{*}\in\{w,w^{\prime},w^{\prime\prime}\}$ such that $d(v_{i},w^{*})\geq 2k+2-i$ for all $11\leq i\leq 2k+$ 1, that is, $b(H(3k+1,3))\geq 2k+2$.
Acknowledgments
The author thanks Naoki Matsumoto for stimulating discussions. The author was supported by JSPS KAKENHI Grant Number JP23K03201.